OpenStudy (anonymous):

Find the exact value by using a half-angle identity. cos 5pi/12

OpenStudy (welshfella):

use the identity cos (x/2) = sqrt[(1 + cos x/ / 2] where cos x = 5pi/6 = - sqrt3/2

OpenStudy (welshfella):

* correction cos (x/2) = sqrt[(1 + cos x) / 2]